# DriveRX: A Vision-Language Reasoning Model for Cross-Task Autonomous Driving

### v1

@@ I.2 Inference Prompt
Since both models are trained on nuScenes-derived data, we conduct a comparison on the DriveBench evaluation set, which also builds upon nuScenes. As shown in Table 11, DriveRX significantly outperforms DriveLMM-o1 on the prediction, planning, and behavior tasks, demonstrating the effectiveness of AutoDriveRL in learning structured, cross-task reasoning. Notably, this result also supports our earlier finding (Section 5.2) that a stronger vision encoder contributes substantially to perception performance. While DriveRX adopts a CLIP-based encoder with efficient visual-language alignment, DriveLMM-o1 leverages the InternVL2.5 encoder, which offers higher capacity for fine-grained visual understanding. As a result, DriveLMM-o1 retains an advantage in the perception task, highlighting the importance of high-resolution visual representations for low-level scene comprehension.
InternVL3 [46] $8$B $33.21$ $41.21$ $70.97$ $51.47$
DriveLMM-o1 [58] $8$B $34.31$ $43.95$ $73.7$ $47.62$
DriveRX $8$B $32.88$ $52.20$ $78.63$ $62.02$

### v2

@@ Appendix I Alignment Between Human and Model Rewards
Since both models are trained on nuScenes-derived data, we conduct a comparison on the DriveBench evaluation set, which also builds upon nuScenes. As shown in Table 15, DriveRX significantly outperforms DriveLMM-o1 on the prediction, planning, and behavior tasks, demonstrating the effectiveness of AutoDriveRL in learning structured, cross-task reasoning. Notably, this result also supports our earlier finding (Section F) that a stronger vision encoder contributes substantially to perception performance. While DriveRX adopts a CLIP-based encoder with efficient visual-language alignment, DriveLMM-o1 leverages the InternVL2.5 encoder, which offers higher capacity for fine-grained visual understanding. As a result, DriveLMM-o1 retains an advantage in the perception task, highlighting the importance of high-resolution visual representations for low-level scene comprehension.
Appendix Q LLM Usage
In this work, large language models (LLMs) were employed as general-purpose assistive tools throughout the research process. Specifically, LLMs were used for text refinement and typo correction to improve the clarity and readability of the manuscript. These contributions were crucial in enhancing the overall quality of the text.
Additionally, LLMs played a significant role in the experimental setup. In our experiments, LLMs were utilized as reward models (RMs) to optimize the performance of reinforcement learning (RL) agents. The LLM-based reward models were trained and evaluated on specific tasks, providing valuable feedback and improving the agent’s decision-making process. This approach was critical in fine-tuning the behavior of the models and achieving desired outcomes in various tasks.
All content generated or refined by LLMs was reviewed and validated by the authors to ensure scientific accuracy and integrity.
Appendix R OOD Evaluation on DriveAction
To further assess the out-of-distribution (OOD) performance of DriveRX, we conducted evaluations on the DriveAction (Hao et al., 2025) benchmark, whose data are sourced from real-world mass-production vehicles with strict manual curation, making it a challenging OOD proxy distinct from DriveRX’s training data.
We employed the official evaluation settings of DriveAction, which utilize an action-rooted tree structure to assess decision-making. Specifically, the benchmark defines four evaluation modes based on the information provided in the prompt to guide the action prediction. First, the Full Pipeline Mode (V-L-A) provides the model with ground-truth QA pairs from both upstream Vision (V) tasks (e.g., traffic light detection) and Language (L) tasks (e.g., navigation following) as textual context, evaluating action performance under fully informed conditions. Second, the Vision/Language-Only Modes (V-A / L-A) inject QA pairs from only one modality into the input to evaluate reliance on specific information. Finally, the Uninformed Mode (A) provides no upstream QA pairs, requiring the model to generate the final action decision relying purely on raw sensor inputs and internal reasoning without high-level external guidance. This mode best reflects the capability of end-to-end autonomous driving systems. We compared DriveRX against baseline models and state-of-the-art (SOTA) VLMs across these metrics, with results presented in Table 16.
The results demonstrate strong OOD robustness. DriveRX significantly outperforms the base model (Align-DS-V) across all metrics on this unseen dataset. Notably, it achieves performance on par with SOTA closed-source models (GPT-4o, o3) and even surpasses most of them on the Action (A) metric (86.18 vs. 81.01/80.67), validating its capability to complex, real-world driving scenarios.
Table 16: OOD Evaluation on DriveAction Benchmark. The results report Action Accuracy.
Model V-L-A V-A L-A A
GPT-4o (OpenAI, 2024a) 88.84 84.72 86.52 81.01
o3 (OpenAI, 2025) 92.19 86.61 88.66 82.23
Claude 3.7 Sonnet (Claude, 2025) 86.31 80.80 82.56 80.67
Align-DS-V (PKU-Alignment, 2025) 78.58 77.32 77.74 76.67
DriveRX 88.09 86.56 87.17 86.18